OpenStudy (anonymous):

series n=0 to infinity ((e^n)+3)/(4^n-1) does it converge and what is its sum?

OpenStudy (anonymous):

\[\int\limits_{0}^{\infty}\frac{ e ^{n}+3 }{ 4^{n}-1 }\]

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

@Kanwar245 Summation,

OpenStudy (anonymous):

oh right

OpenStudy (anonymous):

no the it is 4^(n-1) the 1 is in the exponent

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty}\frac{ e ^{n}+3 }{ 4^{n-1} }\] ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok so each term is positive

OpenStudy (anonymous):

use ratio or root test

OpenStudy (anonymous):

use root test

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} \sqrt[n]{a _{n}}\]

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty}\sqrt[n]{\frac{ e ^{n}+3 }{ 4^{n-1} }}\]

OpenStudy (anonymous):

i separated the the fraction into two fractions and used geometric rule, but having trouble with it.

OpenStudy (anonymous):

okay so when you separate it, on the first part use root test, and the second part is a geometric series of 1/4 which converges

OpenStudy (anonymous):

first part would also converge by root test

OpenStudy (anonymous):

I don't know how would you find the sum of the series though

OpenStudy (anonymous):

for that you would need to look at the partial sums and find their sums

OpenStudy (anonymous):

\[a _{n}=3^{n+1}/2^{n}\]